OpenStudy (anonymous):

Find a cubic function with the given zeros: 7, -3, 2

OpenStudy (anonymous):

options for answer: f(x) = x3 - 6x2 - 13x - 42 f(x) = x3 - 6x2 + 13x + 42 f(x) = x3 - 6x2 - 13x + 42 f(x) = x3 + 6x2 - 13 + 42

OpenStudy (anonymous):

If \(a\) is a root of a polynomial \(P(x)\), then \((x-a)\) divides \(P(x)\) Here we need a polynomial with roots 7, -3, 2 The simplest polynomial that we can think of with these roots is \[(x-7)(x-(-3))(x-2)\] that is \((x-7)(x+3)(x-2)\) Multiply it out to get the result.

OpenStudy (anonymous):

oh okay thanks